Let us consider a quadratic polynomial f(x) such that equation f(x)=7x-14 has exactly one root, and equation f(x)=6-3x has exactly one root. Find the minimum value of a discriminant of f(x).
1
Expert's answer
2020-12-02T01:50:35-0500

Let "f(x)=ax^2+bx+c, a\not=0." Then the discriminant if "f(x)" is



"D=b^2-4ac"

The equation "f(x)=7x-14" has exactly one root



"ax^2+bx+c=7x-14""ax^2+(b-7)x+(c+14)=0""(b-7)^2-4a(c+14)=0"

The equation "f(x)=6-3x" has exactly one root



"ax^2+bx+c=6-3x""ax^2+(b+3)x+(c-6)=0""(b+3)^2-4a(c-6)=0"


We have the system



"b^2-14b+49-4ac-56a=0""b^2+6b+9-4ac+24a=0""b^2-4ac=14b+56a-49""b^2-4ac=-6b-24a-9""14b+56a-49=-6b-24a-9""b=-4a+2""b^2-4ac=24a-12-24a-9""D=-21"



The value of the discriminant of "f(x)" is "-21."